# Sparse Imagination for Efficient Visual World Model Planning

### v1

@@ 4.4 Token Dropout Methods
Type Methods Pointmaze Wall PushT Granular Rope Block Pushing
Random Sampling Random 85.8 93.8 50.4 70.0 73.3 20.5
Fixed 87.1 85.0 47.4 - - 16.5
LHS mckay2000comparison 84.2 92.5 49.6 - - 21.0
Learning-Based Pruning LTRP luo2024ltrp 79.2 94.6 40.9 65.0 42.2 16.5
Attention-Based Pruning Attention-Encoder 79.2 97.1 43.9 68.4 63.3 20.5
STAR Zhang2023star 82.9 85.8 46.9 - 63.3 17.0
Attention-WM 80.0 93.8 44.3 - 61.7 16.5
Cluster & Merging ATC haurum2024agglomerative 78.3 81.2 39.1 25.0 18.3 11.5
Across diverse datasets, our empirical results consistently demonstrate that none of the evaluated token selection methods achieve performance significantly superior to the Random (Table 3). We visualize the actual dropout patterns of different token dropout methods in Figure 7 to illustrate their characteristics. Further experimental results are provided in the Appendix.
@@
Figure 7: Visualization of token dropout patterns for different token reduction methods. Patterns at 50% dropout ratio are shown from left to right for Random, LHS, LTRP, Attention-Encoder, STAR, and Attention-WM.
Spatial Uniformity Does Not Matter.
Comparing LHS (enforcing spatial uniformity) with Random reveals that a purely spatial selection criterion offers no significant advantage in planning success rates. Contrary to an intuition that uniform spatial coverage might be beneficial, our results suggest simply distributing selections evenly is insufficient for prioritizing tokens critical for effective planning.
Random vs. Sophisticated Dropout Methods.
We observe that various token dropout methods do not yield significantly superior performance compared to random dropout. We attribute this to two factors: 1) the substantial inherent redundancy of ViT representations raghu2021vision ; pan2021ia ; chen2022principle ; kim2024learning , and 2) our world model’s training design using randomized grouped attention, which encourages learning to predict from limited token subsets by leveraging this information redundancy.
Random vs. Fixed Dropout Patterns during Planning Steps.
Investigating token reduction strategies, we compared Random and Fixed dropout patterns. Random dropout offers a key robustness advantage in sparse imagination: its stochastic nature enables resampling, facilitating recovery from temporarily dropping essential information in subsequent MPC iterations. This contrasts with fixed patterns, which lack this recovery mechanism. This robustness could explain Random’s empirical outperformance relative to Fixed.
Dropout vs. Clustering-Based Merging.
Lastly, we compare our stochastic dropout (Random) with clustering-based token merging (ATC), which also aims to compress scene representations. Table 3 shows ATC significantly underperforming Random, a gap widening with increased compression. This likely stems from clustering aggregating patches in a manner that diverges from the original distribution and loses fixed spatial positions, hindering world model generalization. Computationally, ATC is significantly slower than Random (up to 268.0% increase planning time), primarily due to the quadratic complexity of clustering and the need for additional spatial matching procedures. Random dropout, conversely, relies on efficient, constant-time index sampling that maintains spatial fidelity. Further details on the clustering implementation and full experimental results for ATC’s planning time are provided in the Appendix.
Information and Probing Analysis of Token Dropout Methods.
HSIC analysis indicates that attention-based selection baselines (e.g., Attention-Encoder) do not achieve significantly higher mutual information than random dropout (Fig. 5). Their often poor nHSIC correlation suggests a misalignment between tokens highlighted by the DINO encoder’s attention and those critical for capturing the state. Also, in attention probing (Fig. 6), compared with the Attention-Encoder baseline, random dropout shows consistently similar performance, questioning the need of additional token selection methods for planning.
@@ 5 Conclusion
In this paper, we introduce a novel approach for computationally efficient planning using sparse imagination, implemented via patch token dropout in transformer-based world models. Our method significantly reduces computational complexity during planning by processing subsets of visual tokens, while incurring only minimal loss in predictive performance. This approach enables fast control in complex visual environments, effectively balancing computational efficiency and task accuracy. Empirical evaluations across various settings confirm the effectiveness and robustness of sparse imagination. Furthermore, our extensive experiments analyze and provide insights into why the framework of sparse imagination with planning is viable. Lastly, our comparative study experimentally demonstrates that diverse token selection strategies do not significantly outperform simple random sampling, suggesting the sufficiency of simple random sampling for planning within this framework.
Limitations
Our work’s key limitation is that experiments were conducted on simulated datasets rather than real-world data. Additionally, although random token dropout is shown effective in reducing computational overhead, the dropout rate requires adjustments based on the specific environment and computational requirements. Finally, our experiments were primarily conducted using a DINO encoder, thus exploring behaviors of other pretrained vision encoders with different representational properties applied to our work is an interesting direction for future research.
@@ Appendix A Implementation Details
Table 4: Dataset details and dataset specific configurations
@@
A.1 Datasets and Environment Details
We elaborate various datasets and environments where our method is experimented on (Table 4). Offline trajectory datasets of Pointmaze, Wall, PushT, Granular and Rope as well as some of dataset-specific configurations are adopted from those provided by [7] [^1] https://github.com/gaoyuezhou/dino_wm. Configurations specific to dataset include history trajectory length for the world model and frame skip length, which is an interval of frames to and from the world model, reducing redundancy between consecutive inputs. We also differ training epoch due to difference of dataset size and complexity.
@@ A.1 Datasets and Environment Details
A navigation task introduced in D4RL [38] where an agent interacts with given 2D maze environment. There are 2000 trajectories with length of 100 created with random actions. Since we use task-agnostic world model with goal-conditioned planning, goal point is not visualized in each frame.
@@
A 2D control task introduced in [39] where an agent needs to manipulate T-shaped block to a designated location. Dataset is generated by rollouting expert demonstration action trajectory with additional action noise in the environment. This allows the world model to learn various dynamics.
@@
Fine-grained manipulation tasks involving deformable objects introduced in [40]. Both datasets have 1000 trajectories with length of 20 created with random actions.

### v2

@@ 4.1 Implementation Details
Pointmaze (60 evals) Wall (60 evals) PushT (60 evals) Granular (20 evals) Rope (30 evals) Block Pushing (50 evals) Avg.
Full 98.3 91.7 75.0 75.0 63.3 22.0 69.8
CLS 96.7 85.0 43.3 20.0 36.7 16.0 50.7
Drop (Ratio)
10% 91.7 93.3 78.3 80.0 73.3 28.0 74.1
20% 90.0 95.0 70.0 80.0 66.7 16.0 69.6
30% 98.3 93.3 61.7 85.0 70.0 18.0 71.0
40% 96.7 93.3 56.7 70.0 76.7 18.0 68.6
50% 100.0 95.0 70.0 60.0 73.3 20.0 69.7
60% 83.3 91.7 46.7 70.0 70.0 20.0 63.6
70% 70.0 93.3 28.3 55.0 53.3 20.0 53.3
80% 73.3 90.0 21.7 50.0 66.7 24.0 54.3
90% 71.7 86.7 20.0 40.0 50.0 12.0 46.7
4.2 Test-Time Trajectory Optimization
Our sparse imagination framework is first evaluated on test-time trajectory optimization using MPC-CEM or CEM. Moderate token dropout (10-50%) maintains task performance relative to the Full-Patch baseline while substantially reducing planning time. For example, in the PushT environment, a 50% dropout ratio cuts the planning time per iteration from 173s to 82s (a 52.6% reduction) without sacrificing performance (Table 1, 2).
This efficiency gain is especially critical in complex environments. Although the CLS-Token offers faster inference, its dramatic performance drop in these scenarios renders it impractical. For instance, in spatially demanding scenes like Granular and Rope, our approach achieves up to 85.0% and 76.7% success respectively, whereas the CLS-Token fails (20.0%, 36.7%) by losing critical spatial information. Similarly, on tasks like PushT, our method with 50% dropout (70.0%) surpasses CLS-Token (43.3%) baselines. Refer to Appendix B.8 for more detailed and fine-grained analysis.
Pointmaze Wall PushT Block Pushing Avg.
Planning Time (s/iter) Change (%) Planning Time (s/iter) Change (%) Planning Time (s/iter) Change (%) Planning Time (s/iter) Change (%) Planning Time (s/iter) Change (%)
Full 184 0.0 79 0.0 173 0.0 297 0.0 183.3 0.0
CLS 49 -73.4 40 -49.4 32 -81.5 163 -45.1 71.0 -61.3
Drop (Ratio)
10% 165 -10.3 73 -7.6 149 -13.9 278 -6.4 166.3 -9.3
20% 141 -23.4 69 -12.7 131 -24.3 259 -12.8 150.0 -18.1
30% 126 -31.5 65 -17.7 114 -34.1 240 -19.2 136.3 -25.6
40% 106 -42.4 62 -21.5 97 -43.9 214 -27.9 119.8 -34.7
50% 93 -49.5 53 -32.9 82 -52.6 208 -30.0 109.0 -40.5
60% 80 -56.5 50 -36.7 69 -60.1 200 -32.7 99.8 -45.6
70% 69 -62.5 46 -41.8 59 -65.9 184 -38.0 89.5 -51.2
80% 56 -69.6 46 -41.8 49 -71.7 175 -41.1 81.5 -55.5
90% 48 -73.9 42 -46.8 38 -78.0 167 -43.8 73.8 -59.8
4.3 Towards Complex and Real-World Tasks: VLA-Guided Planning
To tackle complex and long-horizon tasks, we integrate our planning framework with the latest VLA model such as SmolVLA (Shukor et al., 2025). We evaluate this approach on the LIBERO-10 and Meta-World simulation and the real-world environments using SO-101 robot arm, focusing on two tasks: PickPlace (picking up a block and placing it into a cup) and Drawer (picking up a block, placing it into a drawer, and then closing the drawer). We find that guiding the search with a pre-trained policy is crucial in these demanding settings. For the real-world LeRobot experiments on PickPlace and Drawer, we used a manually selected goal image and proprioceptive state from the dataset for planning (See Appendix A.8 for details).
Simulation Results.
The SmolVLA policy used for simulation experiments is trained to predict action chunks spanning 50 steps and at each decision step we sample 50 candidate action chunks from the policy for LIBERO-10 and 10 for Meta-World to perform planning-based evaluation. In the original work of SmolVLA, only a single action is performed in every policy inference. Since we find it impractial to deploy single-step policies in real world, we execute all actions in the action chunk despite lower absolute performance. In LIBERO-10 evaluation, we test our method in the evaluation benchmark setting with total of 100 trials. Our method (50% token drop, $\bigstar$ in the Fig. 5) boosted the base VLA’s success rate from 29% to 33%. This performance matches the computationally expensive Full-Patch planner but completes episodes nearly twice as fast (29.7s vs. 53.4s per episode). Although it is slower than the base policy due to additional planning overhead, this can be efficiently handled in real-world deployment which will be stated in following section. In Meta-World evaluation, we test on 50 tasks with 15 trials each. As shown in Table 3, our method also show comparable performance to the Full-Patch planner with lower computation overhead.
Table 3: Performance results and planning time across difficulties in Meta-World simulation.
All (%) Easy (%) Medium (%) Hard (%) Very Hard (%) Planning Time (s/episode)
VLA-only 41.87 60.95 20.61 17.78 10.67 1.95
Full 48.80 66.90 29.09 28.89 14.67 3.63
Drop 50% 47.73 65.24 27.27 33.33 12.00 2.37
Real-world LeRobot Results.
In our real-world LeRobot evaluation on PickPlace and Drawer, we tested by fixing the cup’s position or drawer pose and placing the block in initial locations unseen during training for both tasks. We sample 10 distinct candidate chunks from the policy at each decision step instead of 50 for real-world deployment. On a physical robot, sparse imagination with a 50% drop had an even greater impact, significantly increasing the success rate from 60% (VLA-only) to 80% and 70% on PickPlace and Drawer, respectively. Again, this matched the Full-Patch planner’s success rate while reducing planner latency from 19.1s to 10.4s on PickPlace and from 14.0s to 10.6s on Drawer per episode. By leveraging asynchronous inference, this major performance gain across both tasks is achieved with minimal overhead compared to the VLA-only baselines (8.3s and 8.8s), proving our method is a practical solution for real-time robotic planning scenarios.
Figure 5: Trade-off between Inference Time and Performance in LeRobot and LIBERO-10. We show performance of SmolVLA with sparse imagination planning with a particular drop ratio in Left: Real-World PickPlace; Center: Real-World Drawer; Right: LIBERO-10; where ($\bigstar$) represent the operating point we highlight (50% drop), which provides a practical choice between speed and reliability.
4.4 Token Reduction Methods
We compare our random sampling against a broad set of token-selection strategies, grouped into four families as summarized in Table 4: (i) Random Sampling (Random, Fixed, LHS), (ii) Learning-Based Pruning (LTRP), (iii) Attention-Based Pruning (Attention-Encoder, STAR, Attention-WM), and (iv) Cluster & Merging (ATC). All methods share the same world model and planning hyperparameters, and retain the same number of tokens for a fair comparison.
Random Sampling. The Random baseline samples a new subset of patch tokens independently at every planning iteration, while Fixed draws a single subset once for each rollout and reuses it for all rollouts within the episode. We additionally include LHS (Latin Hypercube Sampling; (McKay et al., 2000)), which performs stratified sampling over the 2D patch grid so that retained tokens are approximately uniformly distributed across the image.
Learning-Based Pruning. LTRP (Learning to Rank Patches; (Luo et al., 2024)) trains a ranking network on top of a pre-trained MAE (He et al., 2022) to predict patch importance based on reconstruction sensitivity; at test time, the top-$K$ tokens according to this learned score are kept.
Attention-Based Pruning. Attention-Encoder and STAR (Zhang et al., 2024b) derive token importance from the self-attention maps of a pre-trained DINO encoder, while Attention-WM uses attention maps from our ViT-based world model itself. In all three cases, we rank tokens by their attention-based scores and retain the top-$K$ as the “important” subset.
Cluster & Merging. ATC (Haurum et al., 2024) clusters spatially similar patch tokens (using DINO features) and replaces each cluster with its average, effectively merging redundant patches. For goal-conditioned planning, both the current observation and goal features are clustered, and the distance between predicted futures and goal is computed after aligning clusters via Hungarian matching. A full description of each method and its training configuration is provided in Appendix A.6.
Despite their complexity, empirical results across various datasets consistently show that none of these methods significantly outperform our simple random sampling (Table 4). Fig. 6 visualizes the distinct dropout patterns for each approach.
Table 4: The average planning success rate (%) achieved by various token dropout and merging methods under different drop ratios. Our results suggest that no method achieves significantly superior performance compared to the Random sampling. “Fixed” is excluded for Granular and Rope because they only use CEM, making “Random” and “Fixed” strategies equivalent. For the “Fixed” variant in Granular and Rope tasks, the Avg. is computed by reusing the drop ratios of the corresponding “Random” setting.
Type Methods Pointmaze Wall PushT Granular Rope Block Pushing Avg.
Random Sampling Random 85.8 93.8 50.4 82.5 67.5 20.5 66.7
Fixed 87.1 85.0 47.4 - - 16.5 64.3
LHS (McKay et al., 2000) 84.2 92.5 49.6 82.5 64.2 21.0 65.7
Learning-Based Pruning LTRP (Luo et al., 2024) 79.2 94.6 40.9 75.0 50.8 16.5 59.5
Attention-Based Pruning Attention-Encoder 79.2 97.1 43.9 70.0 67.5 20.5 63.0
STAR (Zhang et al., 2024b) 82.9 85.8 46.9 72.5 63.3 17.0 61.4
Attention-WM 80.0 93.8 44.3 82.5 66.7 16.5 64.0
Cluster & Merging ATC (Haurum et al., 2024) 78.3 81.2 39.1 15.0 25.0 11.5 41.7
@@ 4.4 Token Reduction Methods
Figure 6: Visualization of token dropout patterns for different token reduction methods. Patterns at 50% dropout ratio are shown from left to right for (a) Random, (b) LHS, (c) LTRP, (d) Attention-Encoder, (e) STAR, and (f) Attention-WM.
5 Ablation and Analysis
5.1 Architectural Generalization
To demonstrate broader applicability, we evaluated our sparse imagination framework with MoCo-v3 ViT-S (Chen et al., 2021), MAE ViT-B (He et al., 2022), and DINOv3 ViT-S (Siméoni et al., 2025) in the PointMaze environment (60 episodes). With a 50% token drop, our method matched the MoCo-v3 Full-Patch baseline’s 96.7% success and the DINOv3 Full-Patch baseline’s 98.3% success, while outperforming the MAE Full-Patch baseline (75.0% vs. 68.3%). These results confirm that sparse imagination generalizes effectively across diverse vision encoders.
5.2 Robustness to Token Sparsity during Planning
Training the world model with our randomized grouped attention is crucial for leveraging these sparse token subsets. This strategy significantly reduces prediction errors on token subsets compared to a standard full-attention model (e.g., 0.016 vs. 0.036 of normalized $L_{2}$ error for 50% drop), which directly translates to improved planning performance (Fig. 4). We also find that the CEM planner is robust enough to handle the small remaining prediction noise, based on the experimental results, which is consistent with prior studies (De Boer et al., 2005; Hu and Hu, 2009; Lale et al., 2024). Detailed experiments evaluating CEM’s performance with visual token errors in our framework are in Appendix B.3.
5.3 Information Sufficiency of Sparse Token Subsets
Our analysis confirms that a sparse subset of tokens is sufficient for effective planning. We found that even a small, random fraction of tokens retains substantial ground-truth state information, often more than the global CLS token, as measured by both an information-theoretic analysis using normalized Hilbert-Schmidt Independence Criterion (Gretton et al., 2005), (Fig. 7) and an attentive probing module (Bardes et al., 2024) (Fig. 8). Methodological details are deferred to the Appendix A.4 and A.5.
Figure 7: nHSIC between visual tokens and environment states. nHSIC scores with respect to dropout rate for random dropout compared to Full, CLS, and Attention-Encoder dropout in Granular dataset. High HSIC scores are maintained even under substantial dropout, indicating minimal loss of critical state information.
Figure 8: Attentive probing performance. Validation losses for state prediction with attentive probing compared to Attention-Encoder baseline in Granular dataset. Surprisingly, a single random token shows predictive accuracy comparable (albeit with variance) to the CLS token, indicating state information is distributed across tokens.
5.4 Explaining the Counter-Intuitive Success of Random Sampling
Counter-intuitively, sophisticated importance-based sampling methods failed to show a clear advantage over simple random sampling, often underperforming it (Table 4 vs. Fig. 6).
The “Blind Spot” Problem of Importance-Based Sampling.
We trace this phenomenon to a fundamental “blind spot” problem. Methods based on static importance metrics, by design, learn to ignore regions of the scene deemed unimportant from the static initial and goal images. This proves fatal in dynamic tasks. If an object of interest enters a blind spot, the predicted visual features for different candidate actions become indistinguishable to the planner. This causes the candidate selection process to collapse, as it cannot identify the best action based on visual progress. The failure is starkly demonstrated in a vision-only planning experiment (Wall, 60 episodes, 50% dropout, proprioception distance is not used): the Attention-Encoder method’s success rate collapsed to 21.7%, whereas random sampling, which avoids such blind spots, remained robust at 58.3%. A control experiment using only proprioception confirmed vision’s essential role, achieving just 16.7% success. We provide more detailed analysis in Appendix B.9, and visualization example of this blind-spot failure in Figs. 9 and  20.
(a) Random
(b) Importance-Based (Attention-Encoder)
Figure 9: Visualization Example of the Blind Spot Problem. Both images show 60% token dropout masks in PushT: (a) Random sampling and (b) an importance-based method (Attention-Encoder). In the importance-based case, the blue ball’s path toward any direction to solve the task is almost entirely covered by masked patches, creating a persistent blind spots where the world model cannot observe the object’s movement and thus fails to plan correctly. Random sampling, by contrast, spreads retained tokens more uniformly, making such task-relevant regions less likely to be systematically ignored.
Robustness from Unbiased Coverage and Information Redundancy.
The success of random sampling lies in its unbiased coverage, which is effective precisely because task-relevant information is highly redundant and distributed. We quantify its lack of bias using “cluster retention entropy”, a metric showing random sampling achieves significantly more uniform coverage than Attention-Encoder (2.37 vs. 1.91 bits, higher is more uniform). The viability of this unbiased strategy is confirmed by a critical finding: planning with only the tokens deemed least important by the Attention-Encoder still yielded a strong 63.7% success rate (compared to 79.2% with the most important tokens; PointMaze, 60 episodes, 50% drop). This proves that a random subset is almost certain to contain sufficient information for planning. This finding of distributed information is further validated by our information-theoretic analyses in Section 5.3. More experimental details are presented in Appendix A.7.
Our analyses resolve the paradox of random sampling’s success: its unbiased coverage prevents the fatal “blind spots” that plague static, importance-based methods. This strategy is effective precisely because, as our experiments show, task-relevant information is broadly distributed across the image, ensuring even a random subset likely contains sufficient information for robust planning.
@@ 6 Conclusion
In this paper, we introduced sparse imagination, a novel approach for computationally efficient planning in transformer-based world models. By processing only a subset of visual tokens via random dropout, our method significantly reduces computational complexity while maintaining high task performance across tasks ranging from simple test-time optimizations to complex real-world applications with the latest VLAs. Our comparative analysis showed that various sophisticated sampling methods failed to outperform simple random sampling, and our investigation into the “blind spot” problem suggests this is due to a fundamental mismatch between their static importance metrics and the dynamic nature of planning.
While more advanced, dynamic-aware strategies might exist, they must justify their inherent complexity and computational overhead. For practical applications under resource constraints, we argue that a simple, near-zero-overhead strategy like random dropout provides a powerful and effective baseline, whose practicality may ultimately outweigh the marginal gains of more complex alternatives. The efficiency gains from this approach can then be reallocated to widen the action search or process a longer history of observations. This work establishes random dropout as a simple, robust, and highly practical baseline for deploying visual world models in real-time robotic applications.
Limitations.
Our work has two primary limitations. First, the efficacy of our approach is inherently dependent on the quality of the underlying visual representations from the pre-trained encoder. Second, while we demonstrate that a moderate dropout ratio is a robust default in our benchmarks, we do not claim that any particular fixed ratio is universally optimal, and the ideal sparsity level is likely task-dependent. Although we explore a preliminary uncertainty-aware adaptive mechanism in Appendix B.6, developing a more sophisticated method to adaptively adjust this ratio remains a promising avenue for future research.
Reproducibility Statement.
We strongly encourage the reproducibility of our work. To facilitate this, we provide the maximum possible level of detail regarding our experimental setup in the Appendix A. This includes a comprehensive description of the datasets used, the full training procedures, model architectures, and detailed hyperparameter settings.
Acknowledgements.
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00345809, Research on AI Robustness Against Distribution Shift in Real-World Scenarios; and No. RS-2023-00222663, Center for Optimizing Hyperscale AI Models and Platforms).
Appendix
@@ Appendix A Implementation Details
A.1 Datasets and Environment Details
We elaborate various datasets and environments where our method is experimented on (Table 5). Offline trajectory datasets of Pointmaze, Wall, PushT, Granular and Rope as well as some of dataset-specific configurations are adopted from those provided by Zhou et al. (2024) [^1] https://github.com/gaoyuezhou/dino_wm. Configurations specific to dataset include history trajectory length for the world model and frame skip length, which is an interval of frames to and from the world model, reducing redundancy between consecutive inputs. We also differ training epoch due to difference of dataset size and complexity.
Table 5: Dataset details and dataset-specific configurations. Note that history length, frame skip, and train epochs are for training world model.
@@ A.1 Datasets and Environment Details
[…]
@@
A navigation task introduced in D4RL (Fu et al., 2020) where an agent interacts with given 2D maze environment. There are 2000 trajectories with length of 100 created with random actions. Since we use task-agnostic world model with goal-conditioned planning, goal point is not visualized in each frame.
@@
A 2D control task introduced in Chi et al. (2024) where an agent needs to manipulate T-shaped block to a designated location. Dataset is generated by rollouting expert demonstration action trajectory with additional action noise in the environment. This allows the world model to learn various dynamics.
@@
Fine-grained manipulation tasks involving deformable objects introduced in Zhang et al. (2024a). Both datasets have 1000 trajectories with length of 20 created with random actions.
@@
LIBERO.
The LIBERO benchmark is a multitask robot learning environment featuring a diverse set of objects and tasks. For training, we utilized only the fixed-view images from the LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-10 suites to train both our world model and the SmolVLA policy. This dataset, comprising 1,693 trajectories with 75 to 505 frames each, is publicly available on Hugging Face (available here [^2] https://huggingface.co/datasets/physical-intelligence/libero). We conducted our evaluation exclusively on the LIBERO-10 benchmark, as its challenging long-horizon tasks provide an ideal testbed for validating the efficacy of our planning framework.
Meta-World.
The Meta-World benchmark (Yu et al., 2021) is a comprehensive robotic manipulation suite designed for multi-task and meta-learning. We utilize the MT50 multi-task dataset, which consists of demonstrations for 50 diverse manipulation tasks spanning a wide range of horizons and difficulty levels. The dataset contains 2,500 episodes, each ranging from 14 to 500 frames, and is publicly available on Hugging Face (available here [^3] https://huggingface.co/datasets/lerobot/metaworld_mt50 ). This collection of demonstrations allows our world model and planning framework to learn dynamics across a broad variety of manipulation behaviors.
LeRobot.
We collected real-world datasets for the task “pick the blue block and place it in the cup (PickPlace)” and “put the blue block in the drawer and close it” (Drawer) via teleoperation, using the LeRobot SO-101 leader-follower system with a fixed-view camera. Each dataset comprises 100 trajectories, each containing between 208 to 459 frames. During data collection for PickPlace, the cup’s position remained fixed while the block was placed at 10 distinct initial positions, with 10 trajectories recorded for each. During data collection for Drawer, the drawer’s initial pose remained fixed while the block was placed at 4 distinct initial positions, with 25 trajectories recorded for each. This dataset was then used to train our world model and the SmolVLA policy. For PickPlace task, the trained agent was tested on 5 novel block positions held out from the training set, with the final success rate calculated over 10 total trials (two attempts for each unseen position). For Drawer task, the trained agent was tested on 2 block positions included in the training set, with the final success rate calculates over 10 total trials (five attempts for each position). The SmolVLA policy was loaded from pretrained VLM weights then only the state projector and the action expert are trained when finetuning. The architectural settings follow SmolVLA with 0.45B parameter in the original work. We also leverage asynchronous inference pipeline of SmolVLA. Whenever the remaining action chunk ratio reaches certain threshold, robot client sends the observation from to the policy server and let the remote server compute the next action chunk in advance.
Table 6: Hyperparameters for SmolVLA.
Name Value
Batch size 64
Training steps 50,000(Lerobot), 100,000(LIBERO-10, Meta-World)
Optimizer AdamW
Learning rate 0.0001
Scheduler Cosine
Warmup steps 1,000
Decay steps 30,000
Action chunk size 50
Async inference thres. 0.5